INFRARED SPECTROSCOPY AND PHOTOCHEMISTRY OF NCCN^{+} AND CNCN^{+} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

When a very dilute mixture of NCCN in neon is codeposited at 4.3 K with a beam of neon atoms that have been excited in a microwave discharge, prominent infrared absorptions of the uncharged isomers CNCN and CNNC appear in the resulting solid. The ν_{3} absorption of NCCN^{+} is also very prominent, and the ground-state ν_{1} + ν_{3} absorption of that product is observed for the first time. Photoisomerization to CNCN^{+}, for which the two CN-stretching fundamentals are identified, occurs in the red spectral region, and photodestruction of both NCCN^{+} and CNCN^{+} occurs in the near ultraviolet. The photochemistry of these two cations is correlated with recently published results of experimental and quantum chemical studies of their electronic energy levels.